OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU MEDAL AND FAN please help!!! What is the value of s – 11.5 when s = –4.6?

OpenStudy (anonymous):

-16.1

OpenStudy (texaschic101):

just sub in -4.6 for s in the other equation s - 11.5 = -4.6 - 11.5 = - 16.1

OpenStudy (anonymous):

If you have a PC then you should be able to use a calculator and figure that out :S

OpenStudy (winsteria):

Thank you both so much I'll fan both of you but can you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (winsteria):

Which equation demonstrates the commutative property? A. –(11 + (–4)) = –11 + 4 B. (25 + 7) + (–12) = 25 + (7 + (–12)) C. 6 + (–13) = –13 + 6 D. 12 + (18 + (–6)) = (12 + 18) + (–6)

OpenStudy (winsteria):

opps question marks you might have to reload the question :/ (its a glitch)

OpenStudy (texaschic101):

commutative property states that you can swap numbers and still get the same answer. a + b = b + a so what do you think the answer is ?

OpenStudy (winsteria):

Maybe D idk?

OpenStudy (texaschic101):

not D

OpenStudy (winsteria):

C?

OpenStudy (texaschic101):

yep...its C

OpenStudy (winsteria):

Thank you so much! :)

OpenStudy (texaschic101):

sure thing :)